### v1

@@ 3.3 Visual Concept Discovery and Learning
We propose to discover visual concepts from a large pool of diverse objects, so that the discovered concepts would be generic and generalizable to different domains. Towards this goal, we adopt the Conceptual Captions 3M [26] dataset, which contains three million images and their captions. As illustrated in Figure 2, our overall goal is to identify the key objects and their associated visual concepts from the image captions, with the help of a large language model as the knowledge source. The candidate visual concepts are then checked for their visual discriminativeness, by verifying if a candidate visual concept can be reliably recognized from the corresponding image. Thus, we can select the concepts that can both be recognized by the VLM in the image and be deemed by the LLM as a relevant attribute for the object described in the caption.
@@
As objects in captions often hold specific dependencies, such as serving as the subject or object of an action, we employ Dependency Parsing [5] along with a set of designed rules (see Appendix for details) to retrieve the words and phrases that potentially correspond to objects in the captions (e.g., “king penguin” from a sentence “a group of king penguins walking in the snow”). We then leverage the LLM as an external knowledge base to obtain visual concepts for recognizing the objects. We design a set of prompts (e.g., “What are useful visual features for distinguishing an {object} in a photo?”) to query the LLM and retrieve relevant concepts for the object of interest. We take the union of the concepts discovered for all objects as the preliminary list of generic visual concepts.
We then propose to filter the concepts so that the selected ones are also visually discriminative. Intuitively, a visual concept should be ranked higher when the VLM can recognize it from the image if and only if when the same concept is proposed by the LLM based on the image caption. Otherwise, the concept either is not specific to the object of interest (i.e. can be recognized from the image but not proposed to describe any object mentioned in the caption), or likely to correspond to non-visual concepts (e.g. “popular among kids” and “social birds” in Figure 2). Specifically, given a concept and a list of image-caption pairs, we define two variables $X$ and $Y$, where $X$ corresponds to the image-concept similarity as measured by a VLM, and the $Y$ is a binary indicator on the caption-concept correspondence according to an LLM. A high $X$ value $x$ indicates that the VLM can recognize the concept from the image well, and a higher $Y$ value $y$ indicates the concept is deemed by the LLM as a relevant attribute for the object described in the caption. We adopt the Mutual Information (MI) to determine the agreement between the VLM and the LLM:
@@
We assume that the vision-language pre-training learns powerful encoders for recognizing visual concepts, but visual concept recognition can be further improved by “re-aligning” the image-text interface. As illustrated in Figure 3, we keep the weights of pre-trained text and image encoders frozen and only fine-tune the last linear projection layers used to compute the concept activations $\mathbf{a}$. We re-purpose the concept-based object recognition framework $f(\mathbf{a})=\mathbf{a}\cdot\mathcal{W}$: Instead of learning the $N\times M$ weight matrix $\mathcal{W}$, we construct a fixed concept-category association matrix $\mathcal{W}_{\text{LLM}}$ with binary weights, where each weight is obtained by querying the LLM with prompts like “Does the {category_name} usually have the attribute {concept}?”. We then fine-tune only the linear projection layers of the VLM through the concept bottleneck, by asking it to justify its own prediction: Namely, the classification objective is obtained by asking the VLM to perform zero-shot object classification. We refer to our two-stage pipeline as the Concept Discovery and Learning (CDL) framework, it leverages knowledge already encoded in pre-trained LLMs and VLMs, and the whole framework is self-supervised as it requires no additional human annotations.
@@ 3.4 Visual Concept Applications
Our concept discovery and learning framework aims to obtain a list of generic and visually discriminative concepts. For object recognition applications, the downstream dataset usually focuses on a specific domain, and it is often desirable to obtain a compact and performant set of visual concepts (e.g. “four-wheeled” is useful for recognizing cars, but not so much for birds). Towards this goal, we first construct the concept-category association matrix $\mathcal{W}_{\text{LLM}}$ and perform concept learning only for the object categories in the target dataset, so that irrelevant concepts not used by any objects are automatically discarded.
@@
We also expect the selected concepts to generalize to unseen but related object categories. We employ a heurstics where a visual concept is more likely to generalize if it is already used by many known object categories. The “generalizability” of a concept $G(c)$ can hence be estimated by the ratio of object categories that contain the concept $c$ over the total number of object categories.
@@ 3.5 Visual Concept Evaluation
Interpretability metric from [39] quantitatively measures how well a VLM learns concept-category associations. The discovered concepts are interpretable if they are associated with images in a human-understandable manner. For example, we expect the model to recognize an image of a “Giant Panda” according to the concepts “Rounded head”, “Black eye patches” and “Black and white fur” instead of irrelevant concepts like “long wings”. This metric is computed with an intervention approach. Given a trained concept-association matrix, we replace the concept activations with the binary ground-truth concept values to make predictions. High intervention accuracy indicates high interpretability of visual recognition since the concept-category association is consistent with human knowledge.
Precision is measured with human evaluation on how well the discovered concepts provide correct reasoning clues for visual recognition. As illustrated in Section 3.1, the contribution of a certain concept to the predicted category $j$ can be measured by the dot product between the concept activation $\mathbf{a}$ and $\mathcal{W}_{j}$ from a CBM weight matrix $\mathcal{W}$. We expect the top-weighted concepts to be accurate depictions of the predicted object when the model prediction is correct. For each correctly-classified image, we ask human annotators to determine whether the top-$k$ weighted concepts actually describe the image, and calculate the proportion as the precision score. The selection of the parameter $k$ is contingent upon the scale of the concept space within each dataset. For datasets with large amounts of concepts, we can find more descriptive concepts for a certain category. Hence, we set a larger value for $k$. Conversely, datasets with relatively few concepts have a diminished pool of descriptive elements for individual categories. Consequently, a smaller value for $k$ is deemed appropriate in such scenarios.
@@ 4.1 Experimental Setup
Datasets: We conduct experiments on several challenging fine-grained image classification datasets, including ImageNet [6], Food-101 [3], CIFAR-100 [11], CIFAR-10, CUB-200 [30] and Flowers-102 [23]. The statistics and split of the datasets are shown in the appendix.
Baselines: We compare with LaBo [35] and LM4CV [32], which are the state-of-the-art works on concept-based visual recognition. Following the setting in LM4CV, we control the bottleneck sizes (number of concepts) to be the same for the baselines and our model for fair comparison. Besides full-shot classification, we also compare with LaBo on few-shot settings, where we select concepts and train the model on limited data. It is infeasible for LM4CV to conduct few-shot classification since it requires the whole training set on concept generation.
Implementation Details: We use the same LLM GPT-3-text-davinci-002 to obtain descriptors as previous works. We also use the same CLIP backbone (ViT-L-14) to compare with baseline models. Following [39], we use logistic regression to train the concept bottleneck models. We observe that the performance of CBM is robust to the choice of hyperparameters and use the default Scikit-learn hyperparameters for all datasets. For concept learning, we use the AdamW optimizer with 5e-4 learning rate and 1e-4 weight decay to fine-tune the CLIP model, and we use the validation loss to select checkpoints. For human evaluation experiments, we hire annotators from Amazon Mechanical Turk. For each example, we ask three annotators to annotate and use the majority vote to obtain the result. We conduct Students’ T-test [27] to evaluate the statistical significance of the human evaluation results. The detailed design of human evaluation and the statistical significance results are shown in the appendix.
@@ 4.2 Concept-Augmented Prompts Are Category-Biased
As discussed in Section 3.2, we conduct ablation experiments to understand if the concepts used in the prompts of previous zero-shot classification methods lead to the improved classification accuracy. The results in Table 1 are consistent with Figure 1, both of which show that the concept-augmented text prompts do not offer conclusive evidence whether pre-trained VLMs learn to encode concepts. In contrast, our proposed CDL provides concept prompts independent of category names, and still achieves competitive zero-shot classification accuracy.
Method Prompt Design ImageNet Food-101 CIFAR-100 CIFAR-10 CUB-200 Flowers-102 VDES Category Name 71.6 91.8 75.9 96.2 63.1 77.4 VDES Name w/ LLM Concepts 75.0 92.4 77.7 96.6 63.5 78.9 VDES Name w/ Random Concepts 70.1 91.6 75.4 95.0 62.1 78.7 VDES LLM Concepts 22.1 3.6 30.9 70.7 5.3 7.0 CDL CDL Concepts 75.7 91.5 77.8 96.5 64.7 80.9
Table 1: Zero-shot classification on six object recognition benchmarks. We observe that although augmenting category names with LLM-generated concepts improves classification accuracy (e.g. VDES [21]), the method still mainly relies on category names in the text prompts, and it remains unclear if the concepts included in the text prompts are properly recognized by pre-trained VLMs. When paired with randomly shuffled concept in the prompts, the accuracy drops only moderately; when the category names are removed, the accuracy drops significantly. Our proposed concept discovery and learning (CDL) framework achieves competitive zero-shot classification performance without using category names as a shortcut.
@@
Table 2: Human evaluation results on the proportion of category-agnostic and visually-discriminative concepts. The results show that our concepts discovered from general corpora are less biased by category names and more visually discriminative.
ImageNet Food-101 CIFAR-100 CIFAR-10 CUB-200 Flowers-102 #Concepts 1000 2000 101 202 100 200 10 20 200 400 102 204 LaBo 83.2 83.6 89.8 91.1 80.5 84.1 77.8 92.2 79.7 81.3 95.5 94.9 LM4CV 83.4 83.6 90.0 91.1 81.7 84.2 80.4 92.3 77.5 80.4 94.5 93.8 CDL 83.6 84.0 94.2 94.9 83.8 85.8 82.1 93.5 83.2 83.4 96.3 95.7
@@
Figure 4: Few-shot classification evaluation with LaBo and our proposed CDL.
@@ 4.3 Classification Performance with Visual Concepts
We then adopt the few-shot learning setting, where the concepts are selected only with the few training examples. Figure 4 shows that CDL consistently outperforms LaBo, especially when the number of training examples is smaller.
@@ 4.4 Evaluation of the Discovered Concepts
Figure 5: Human evaluation of the discovered concepts for their interpretability, precision, and thoroughness.
@@
Figure 5 shows the evaluation results of the Interpretability, Precision, and Thoroughness of the baselines and our method. We can observe that despite having high classification performance, the baseline models discover concepts that exhibit unsatisfactory interpretability, precision, and thoroughness. In contrast, our CDL framework provides significant improvements on all three metrics.
@@
$\mathrm{\Delta}$Classification $\mathrm{\Delta}$Interpretability $\mathrm{\Delta}$Precision $\mathrm{\Delta}$Thoroughness LaBo 0.2 16.1 5.8 7.6 LM4CV 0.4 14.2 4.5 6.4 CDL 1.0 18.3 7.9 10.4
@@
Table 4: In-domain generalization evaluation, where $\mathrm{\Delta}$ denotes the improvement of fine-tuned CLIP compared to the original CLIP on the unseen categories.
$\mathrm{\Delta}$Classification $\mathrm{\Delta}$Interpretability $\mathrm{\Delta}$Precision $\mathrm{\Delta}$Thoroughness LaBo -1.5 0.1 -3.7 -4.5 LM4CV -0.3 -0.6 -6.4 -8.2 CDL 0.2 7.0 10.3 13.9
Table 5: Cross-domain generalization evaluation from ImageNet to CUB-200.
@@
Tables 4 and 5 show the in-domain and cross-domain generalization results. , and Figure A3 shows the generalization of our discovered concepts under the few-shot setup. Our proposed CDL outperforms both baselines for its generalizability, especially for the cross-domain scenario when transferring from ImageNet to CUB-200. We observe that LaBo and LM4CV struggle with cross-domain generalization as they select completely different concepts for different datasets and few common patterns can be learned with their methods.
Table 6: Human evaluation of the discovered concepts before and after concept learning on the ImageNet dataset.
Finally, Table 6 shows the human evaluation results before and after concept learning on the ImageNet dataset. We can observe both our concept discovery and concept learning can improve the quality of the extracted concepts compared with baselines.

### v2

@@ 3.3 Visual Concept Discovery and Learning
We propose to discover visual concepts from a large pool of diverse objects, so that the discovered concepts would be shared by multiple objects and generalizable to different domains. Towards this goal, we adopt the Conceptual Captions 3M (Sharma et al., 2018) dataset (CC-3M). CC-3M contains three million images and their captions, hence offering a huge repertoire of objects and their characteristics rendered in both images and text. As illustrated in Figure 2, our overall goal is to identify the key objects and their associated visual concepts from the image captions, with the help of a large language model as the knowledge source. The candidate visual concepts are then checked for their visual discriminativeness, by verifying if a candidate visual concept can be reliably recognized from the corresponding image. Thus, we can select the concepts that can both be recognized by the VLM in the image and be deemed by the LLM as a relevant attribute for the object described in the caption.
@@
As objects in captions often hold specific dependencies, such as serving as the subject or object of an action, we employ Dependency Parsing (De Marneffe et al., 2006) along with a set of designed rules (see Appendix for details) to retrieve the words and phrases that potentially correspond to objects in the captions (e.g., “king penguin” from a sentence “a group of king penguins walking in the snow”). We then leverage the LLM as an external knowledge base to obtain visual concepts for recognizing the objects. We design a set of prompts (e.g., “What are useful visual features for distinguishing an {object} in a photo?”) to query the LLM and retrieve relevant concepts for the object of interest. We take the union of the concepts discovered for all objects as the preliminary list of generic visual concepts.
We then propose to filter the concepts so that the selected ones are also visually discriminative. Intuitively, a visual concept should be ranked higher when the VLM can recognize it from the image if and only if when the same concept is proposed by the LLM based on the image caption. Otherwise, the concept either is not specific to the object of interest (i.e. can be recognized from the image but not proposed to describe any object mentioned in the caption), or likely to correspond to non-visual concepts (e.g. “popular among kids” and “social birds” in Figure 2). Specifically, given a concept and a list of image-caption pairs, we define two variables $X$ and $Y$, where $X$ corresponds to the image-concept similarity as measured by a VLM, and the $Y$ is a binary indicator on the caption-concept correspondence according to an LLM. A high $X$ value $x$ indicates that the VLM can recognize the concept from the image well, and a higher $Y$ value $y$ indicates the concept is deemed by the LLM as a relevant attribute for the object described in the caption. In practice, we compute $x$ as the cosine similarity between the CLIP embedding of the concept and the image, and define $y$ as a boolean indicating whether the concept is relevant to the object in the image, as determined by the LLM. We adopt the Mutual Information (MI) to determine the agreement between the VLM and the LLM:
@@
Although we can discover general and visually discriminative concepts from pre-trained VLMs, these concepts are not always perfectly-align with images as shown in Column 2 of Figure 1 and Row 4 of Table 1. We propose a self-supervised fine-tuning method to refine the image-concept alignment of VLMs by leveraging the knowledge already encoded in pre-trained VLMs and LLMs. We first perform the zero-shot classification with pre-trained VLMs to generate pseudo labels for training. We then construct a “ground-truth” concept-category association matrix $\mathcal{W}_{\text{LLM}}$ with binary weights, where each weight is obtained by querying the LLM with prompts like “Does the {category_name} usually have the attribute {concept}?”. As illustrated in Figure  3, we project the image and concept embeddings into concept activations $\mathbf{a}$ using the linear projection. The prediction of image category is obtained by the matrix multiplication, where $f(\mathbf{a})=\mathbf{a}\cdot\mathcal{W}_{\text{LLM}}$. Assuming that the $\mathcal{W}_{\text{LLM}}$ provides ground-truth concept-category association, the loss derived from the predicted labels is utilized to optimize the image-concept alignment. The whole framework is self-supervised as it requires no additional human annotations.
@@ 3.4 Visual Concept Applications
Our concept discovery and learning framework aims to obtain a list of generic and visually discriminative concepts. We propose to apply the discovered concepts on fine-grained object recognition benchmarks such as bird classification (Wah et al., 2011) and food classification (Bossard et al., 2014) to evaluate their quality. For such applications, it is often desirable to obtain a compact and performant set of visual concepts (e.g. “four-wheeled” is useful for recognizing cars, but not so much for birds). Towards this goal, we first construct the concept-category association matrix $\mathcal{W}_{\text{LLM}}$ and perform concept learning only for the object categories in the target dataset, so that irrelevant concepts not used by any objects are automatically discarded.
@@
We also expect the selected concepts to be generalizable, that is, to benefit the recognition of unseen objects. We employ a heuristics where a visual concept is more likely to generalize if it is already used by many known object categories. The “generalizability” of a concept $G(c)$ can hence be estimated by the ratio of object categories that contain the concept $c$ over the total number of object categories.
@@ 3.5 Visual Concept Evaluation
Interpretability metric from Yun et al. (2023) quantitatively measures how well a VLM learns concept-category associations. The discovered concepts are interpretable if they are associated with images in a human-understandable manner. For example, we expect the model to recognize an image of a “Giant Panda” according to the concepts “Rounded head”, “Black eye patches” and “Black and white fur” instead of irrelevant concepts like “long wings”. In this way, the discovered concepts should lead to a concept-category association matrix that is similar to $\mathcal{W}_{LLM}$. Given a trained concept-category matrix $\mathcal{W}_{trained}$, we can predict the category for an image with its concept activations, that is, $f(\mathbf{a})=\mathbf{a}\cdot\mathcal{W}_{trained}$. Following Koh et al. (2020); Yun et al. (2023), we intervene the classification with ground-truth concept activations $\mathbf{g}$, where $\mathbf{g}_{i}=1$ if the ground-truth category of the image contains the $i$-th concept according to the LLM knowledge and $\mathbf{g}_{i}=0$ otherwise. We calculate the accuracy of predictions made by $f(\mathbf{g})=\mathbf{g}\cdot\mathcal{W}_{trained}$ as the intervention accuracy. High intervention accuracy indicates high interpretability of visual recognition since the concept-category association is consistent with human knowledge.
Precision is measured with human evaluation on how well the discovered concepts provide correct reasoning clues for visual recognition. As illustrated in Section 3.1, the contribution of a certain concept to the predicted category $j$ of an image can be measured by the dot product between the concept activation $\mathbf{a}$ and $\mathcal{W}_{j}$ from a CBM weight matrix $\mathcal{W}$. We expect the top-weighted concepts to be accurate depictions of the object of interest in the image when the model prediction is correct. For each correctly-classified image, we ask human annotators to determine whether the top-$k$ weighted concepts actually describe the image and calculate the proportion as the precision score. The selection of the parameter $k$ is contingent upon the scale of the concept space within each dataset. For datasets with large amounts of concepts, we can find more descriptive concepts for a certain category. Hence, we set a larger value for $k$. Conversely, datasets with relatively few concepts have a diminished pool of descriptive elements for individual categories. Consequently, a smaller value for $k$ is deemed appropriate in such scenarios.
@@ 4 Experiments
We conduct quantitative and human evaluations to evaluate the discovered and learned concepts. In Section 4.2, we demonstrate why we need short-cut free and visually discriminative concept discovery and learning by showing that concept prompts in previous works might be category-biased. In Section 4.3 and Section 4.4, we aim to prove that pre-trained VLMs do learn visual concepts by demonstrating that the discovered concepts can lead to accurate classification and exhibit high qualities including interpretability, precision, thoroughness and generalizability. In Section 4.5 we aim to explore how specific components in our proposed framework contribute to the quality of discovered concepts.
@@ 4.1 Experimental Setup
Datasets: We conduct experiments on several challenging fine-grained image classification datasets, including ImageNet (Deng et al., 2009), Food (Bossard et al., 2014), CIFAR-100 (Krizhevsky et al., 2009), CIFAR-10, CUB (Wah et al., 2011) ,Flowers (Nilsback & Zisserman, 2008) Stanford Cars (Krause et al., 2015), Aircrafts (Maji et al., 2013) and Oxford Pets (Parkhi et al., 2012). The statistics and split of the datasets are shown in the appendix.
Baselines: We compare with LaBo (Yang et al., 2023b) and LM4CV (Yan et al., 2023a), which are the state-of-the-art works on concept-based visual recognition. Following the setting in LM4CV, we control the bottleneck sizes (number of concepts) to be the same for the baselines and our model for fair comparison. Besides full-shot classification, we also compare with LaBo on few-shot settings, where we select concepts and train the model on limited data. It is infeasible for LM4CV to conduct few-shot classification since it requires the whole training set on concept generation.
Implementation Details: We use the same LLM GPT-3-text-davinci-002 to obtain descriptors as previous works. We also use the same CLIP backbone (ViT-L-14) to compare with baseline models. Following (Yun et al., 2023), we use logistic regression to train the concept bottleneck models. We observe that the performance of CBM is robust to the choice of hyperparameters and use the default Scikit-learn hyperparameters for all datasets. For concept learning, we use the AdamW optimizer with 5e-4 learning rate and 1e-4 weight decay to fine-tune the CLIP model, and we use the validation loss to select checkpoints. For human evaluation experiments, we hire annotators from Amazon Mechanical Turk. For each example, we ask three annotators to annotate and use the majority vote to obtain the result. We conduct Students’ T-test (Student, 1908) to evaluate the statistical significance of the human evaluation results. The detailed design of human evaluation and the statistical significance results are shown in the appendix.
@@ 4.2 Concept-Augmented Prompts Are Category-Biased
As discussed in Section 3.2, we conduct ablation experiments to understand if the concepts used in the prompts of previous zero-shot classification methods lead to improved classification accuracy. The results in Table 1 are consistent with Figure 1, both of which show that the concept-augmented text prompts do not offer conclusive evidence whether pre-trained VLMs learn to encode concepts. Even random concepts have minimal impact on zero-shot performance (as shown in Row 2 and Row 3). In contrast, the removal of category names results in a catastrophic decline in zero-shot performance (as shown in Row 2 and Row 4). These observations suggest that category names act as critical shortcuts in concept-augmented prompts.
Prompt Design ImageNet Food CIFAR-100 CIFAR-10 CUB Flowers Stanford Cars Aircrafts Oxford Pets Category Name 71.6 91.8 75.9 96.2 63.1 77.4 77.3 36.1 93.5 Name w/ LLM Concepts (Menon & Vondrick, 2022) 75.0 92.4 77.7 96.6 63.5 78.9 77.6 37.4 92.2 Name w/ Random Concepts 70.1 91.6 75.4 95.0 62.1 78.7 76.5 36.0 92.4 LLM Concepts only 22.1 3.6 30.9 70.7 5.3 7.0 3.2 1.6 6.4
Table 1: Zero-shot classification on nine object recognition benchmarks. We observe that although augmenting category names with LLM-generated concepts improves classification accuracy (e.g. VDES (Menon & Vondrick, 2022)), the method still mainly relies on category names in the text prompts, and it remains unclear if the concepts included in the text prompts are properly recognized by pre-trained VLMs. When paired with randomly shuffled concept in the prompts, the accuracy drops only moderately; when the category names are removed, the accuracy drops significantly.
@@
Table 2: Human evaluation results on the proportion of category-agnostic and visually-discriminative concepts. The results are averaged over all datasets. The results show that our concepts discovered from general corpora are less biased by category names and more visually discriminative.
ImageNet Food CIFAR-100 CIFAR-10 CUB Flowers Stanford Cars Aircrafts Oxford Pets #Concepts 1000 2000 101 202 100 200 10 20 200 400 102 204 196 392 102 204 37 74 LaBo 83.2 83.6 89.8 91.1 80.5 84.1 77.8 92.2 79.7 81.3 95.5 94.9 84.7 85.6 59.3 60.7 88.3 90.7 LM4CV 83.4 83.6 90.0 91.1 81.7 84.2 80.4 92.3 77.5 80.4 94.5 93.8 84.9 85.7 59.9 61.0 88.5 90.4 CDL 83.6 84.0 94.2 94.9 83.8 85.8 82.1 93.5 83.2 83.4 96.3 95.7 86.1 86.7 60.8 62.4 90.6 92.3
@@
Figure 4: Few-shot classification evaluation with LaBo and our method.
@@ 4.3 Classification Performance with Visual Concepts
We then adopt the few-shot learning setting, where we select the concepts (as described in 3.4) and train the model with the few training examples. Figure 4 shows that CDL consistently outperforms LaBo, especially when the number of training examples is smaller.
@@ 4.4 Evaluation of the Discovered Concepts
Figure 5: Human evaluation of the discovered concepts for their interpretability, precision, and thoroughness. We evaluate the concepts selected without and with concept learning (denoted as “CD” and “CDL”).
@@
Figure 5 shows the evaluation results of the Interpretability, Precision, and Thoroughness of the baselines and our method. For human evaluation of Precision, and Thoroughness, we randomly select 100 images from each dataset and ask 3 human workers to annotate the Precision and Thoroughness of top-weighted concepts for each image as described in 3.5. More details are shown in Appendix Sec. (D). The Precision and Thoroughness of each dataset are calculated as the averages of these evaluations on the 100 images, with the results displayed on the Y-axis of Figure 5. We can observe that despite having high classification performance, the baseline models discover concepts that exhibit unsatisfactory interpretability, precision, and thoroughness. Our concept discovery method provides significant improvements on all three metrics, and the self-supervised concept learning can further improve the concept quality.
@@
Domain Method $\mathrm{\Delta}$Classification $\mathrm{\Delta}$Interpretability $\mathrm{\Delta}$Precision $\mathrm{\Delta}$Thoroughness In-domain LaBo 0.2 16.1 5.8 7.6 LM4CV 0.4 14.2 4.5 6.4 CDL 1.0 18.3 7.9 10.4 Cross-domain LaBo -1.5 0.1 -3.7 -4.5 LM4CV -0.3 -0.6 -6.4 -8.2 CDL 0.2 7.0 10.3 13.9
@@
Table 4: Generalization evaluation for discovered concepts. The In-domain results refer to improvements on unseen categories in the same dataset, while the Cross-domain results refer to improvements on unseen domains. $\mathrm{\Delta}$ denotes the improvement of fine-tuned CLIP compared to the original CLIP. Higher improvement means better generalization.
@@
Table 4 shows the in-domain and cross-domain generalization results. Our proposed CDL outperforms both baselines for its generalizability, especially for the cross-domain scenario when transferring from ImageNet to CUB. We observe that LaBo and LM4CV struggle with cross-domain generalization as they select completely different concepts for different datasets and few common patterns can be learned with their methods.
4.5 Ablation Study
In this subsection, we perform ablation studies to analyze the performance of different sub-modules in the proposed CDL framework.
Multi-modal Visual Concept Discovery
To evaluate the effectiveness of the proposed concept discovery method, we build a baseline which only uses the LLM to generate concepts and perform random concept selection for downstream tasks. As shown in Table 5, the concepts discovered by the LLM-only method are of unsatisfactory classification accuracy, interpretability, precision and thoroughness, which demonstrates the necessity of utilizing multi-modal information to select visually discriminative concepts. The results show that our proposed multi-modal concept discovery method can effectively improve the classification performance and the interpretability, precision and thoroughness of the discovered concepts.
Accuracy Interpretability Precision Thoroughness LLM-only 68.5 3.6 24.2 27.0 Multi-modal 83.4 44.6 67.3 71.0
Table 5: Comparison of our proposed multi-modal concept discovery method with the LLM-only concept discovery method on CUB dataset. Our method significantly outperforms LLM-only method on both classification accuracy and concept quality.
Utilization of the CC-3M dataset
To analyze the effectiveness of the utilization of the CC-3M dataset during concept discovery, we perform the proposed concept discovery method on the CC-3M (Sharma et al., 2018) dataset and the downstream datasets. The performances of both methods are similar on CUB (details are shown in Appendix), which indicates that our proposed concept discovery method can still provide significant improvement on the classification and concept quality with only downstream datasets. Considering the cost of human evaluation for the Precision and Thoroughness, we select the CUB dataset as the downstream benchmark for the generalization evaluation. The in-domain and cross-domain generalization results in Table 6 indicate that the concepts discovered from CC-3M are more generalizable than the concepts discovered from downstream datasets. In conclusion, the proposed concept discovery and learning method provides improvement on classification performance and concept quality regardless of the utilization of datasets, and discovering concepts from the CC-3M datasets can enhance the generalization of concepts.
Domain Source $\mathrm{\Delta}$Classification $\mathrm{\Delta}$Interpretability $\mathrm{\Delta}$Precision $\mathrm{\Delta}$Thoroughness In-domain CUB 0.2 12.5 2.8 3.7 CC-3M 1.0 18.3 7.9 10.4 Cross-domain ImageNet -1.1 -0.2 -5.5 -4.6 CC-3M 0.2 7.0 10.3 13.9
Table 6: Comparison of the concept discovered from CC-3M and other datasets on both in-domain and cross-domain generalization evaluation for the CUB task. The concepts discovered from CC-3M have better generalizability in both cases.